# Symbolic Feedforward Networks for Probabilistic Finite Automata: Exact Simulation and Learnability

### v1

@@ Abstract
We present a formal and constructive theory showing that probabilistic finite automata (PFAs) can be exactly simulated using symbolic feedforward neural networks. Our architecture represents state distributions as vectors and transitions as stochastic matrices, enabling probabilistic state propagation via matrix-vector products. This yields a parallel, interpretable, and differentiable simulation of PFA dynamics using soft updates—without recurrence. We formally characterize probabilistic subset construction, $\varepsilon$-closure, and exact simulation via layered symbolic computation, and prove equivalence between PFAs and specific classes of neural networks. We further show that these symbolic simulators are not only expressive but learnable: trained with standard gradient descent-based optimization on labeled sequence data, they recover the exact behavior of ground-truth PFAs. This learnability, formalized in Proposition 5.1, is the crux of this work. Our results unify probabilistic automata theory with neural architectures under a rigorous algebraic framework, bridging the gap between symbolic computation and deep learning.
@@ 1 Introduction
This work closes that gap. We show that feedforward networks—without recurrence, memory, or sampling—can simulate PFAs exactly. Each layer in our architecture implements a stochastic state update using matrix-vector multiplication with symbolic transition matrices, followed by a softmax or sigmoid operation for normalization and classification. Our framework extends prior work on DFA Dhayalkar (2025c) and NFA Dhayalkar (2025d) simulation by unifying deterministic, nondeterministic, and probabilistic transitions under a single symbolic neural model.
Formally, we define the simulation of PFAs by feedforward networks as the exact computation of the acceptance probability for any string via symbolic matrix composition. Proposition 4.7 and Theorem 4.10 prove a bidirectional equivalence: (1) every PFA can be simulated by a depth-unrolled feedforward network with shared stochastic parameters, and (2) every such symbolic network corresponds to a valid PFA. This allows us to replace automaton-based reasoning with static, interpretable network computation. Moreover, we define new symbolic mechanisms for probabilisitc $\varepsilon$-closure (Lemma 4.6), probabilisitc subset construction (Theorem 4.5), and belief state propagation (Proposition 4.1), all implemented as differentiable neural layers.
@@
  • We present the first constructive theory for simulating probabilistic finite automata (PFAs) using symbolic feedforward networks.
  • We introduce exact formulations of probabilistic subset construction and $\varepsilon$-closure via matrix-vector dynamics, enabling differentiable automaton emulation.
@@
This work advances the field of neural-symbolic computation by providing a unified, interpretable, and learnable framework for simulating probabilistic automata in neural networks. It lays the groundwork for deeper integration of formal language theory and deep learning, with applications in structured reasoning, interpretable sequence modeling, and symbolic control. This paper uses a structure with lemma, propositions, remarks, theorems and corollaries, similar to prior work Zhang et al. (2015); Hanin and Sellke (2018); Neyshabur et al. (2015); Dhayalkar (2025d; b; a).
@@ 2 Related Work
The study of neural networks as computational models has a long and rich history, spanning universal approximation theory Cybenko (1989); Hornik et al. (1989), neural-symbolic integration Besold et al. (2017); d’Avila Garcez et al. (2019), automata learning Weiss et al. (2018); Hupkes et al. (2020); Michalenko (2019), and formal language recognition Hopcroft et al. (2006); Hays (2003). This paper builds on this foundational work by developing a precise theoretical and empirical framework that characterizes how feedforward networks simulate probabilistic finite automata (PFAs), extending prior results from the deterministic setting Dhayalkar (2025c) and non-deterministic setting Dhayalkar (2025d).
@@
Neural-Symbolic Computation. Early research in neural-symbolic computation sought to integrate symbolic reasoning with neural networks via hybrid architectures or recurrent logic units d’Avila Garcez et al. (2019); Besold et al. (2017). These models demonstrated that certain forms of propositional logic could be encoded and manipulated within connectionist systems, yet often lacked rigorous formalism and focused primarily on logic inference rather than state-machine emulation. More recent efforts have introduced differentiable models of symbolic processes Graves (2017), neural theorem proving Rocktäschel and Riedel (2017), and neural architecture search over structured domains Łukasz Kaiser et al. (2017). However, these methods have been largely heuristic or data-driven, often requiring supervision or pre-specification of symbolic structure.
@@
Neural Networks as Automata. Multiple works have explored whether and how neural networks implicitly implement automata when trained on regular or formal languages Weiss et al. (2018); Hupkes et al. (2020); Rabinovich et al. (2017). These studies generally relied on probing internal representations of recurrent networks (RNNs, LSTMs, Transformers) to detect automaton-like behaviors such as state memorization or finite counting. Yet, such work remained empirical in nature, offering limited constructive insight into network design or simulation guarantees. Our work, building upon Dhayalkar (2025d), replaces this probing-based approach with a complete, formal construction that shows how feedforward networks can simulate finite automata exactly—transition by transition—both in deterministic and probabilistic settings. We go beyond approximate modeling to provide exact emulation and generalization guarantees.
@@
Formal Automata Theory. The classical theory of automata Hopcroft et al. (2006); Hays (2003) distinguishes deterministic, nondeterministic, and probabilistic automata (DFAs, NFAs, PFAs), with growing expressive power but also increasing simulation complexity. In the deterministic case, transitions are deterministic functions, while in PFAs, transitions are stochastic matrices operating on probability distributions over states. While PFAs have been studied extensively in formal language theory Rabin (1963); Paz (1971), their neural simulation has received little attention. Most existing neural-symbolic models either avoid stochastic transitions or simulate them via recurrent dynamics, often without exactness guarantees. We fill this gap by providing the first constructive simulation of PFAs using feedforward neural networks, showing that symbolic probabilistic state updates, $\varepsilon$-closures, and subset constructions can all be implemented using static, differentiable architectures.
@@
Probabilistic and Nondeterministic Extensions. Recent advances in learning finite automata machines Siegelmann (1996); Giles et al. (1992b); Hong et al. (2022); Omlin and Giles (1996); Marzen and Crutchfield (2022); Wei et al. (2022) explore how RNNs approximate NFAs and pushdown automata, yet lack constructive bounds or clarity about how transitions or path probabilities are represented. Prior work by  Dhayalkar (2025c) introduced a constructive theory for simulating deterministic finite automata (DFAs) using depth-unrolled feedforward neural networks with ReLU or threshold activations. That work demonstrated exact transition emulation, formally proved state transition separability and embedding capacity, and connected network representations to the classical Myhill-Nerode theorem. Subsequently, this framework was extended to the nondeterministic case Dhayalkar (2025d), showing that feedforward architectures could simulate nondeterministic finite automata (NFAs) without requiring explicit path enumeration. The NFA variant introduced symbolic binary state vectors and unrolled subset constructions to represent all reachable states at each timestep. Importantly, it preserved full theoretical interpretability while maintaining exact correspondence with NFA behavior, including support for $\varepsilon$-transitions.
@@
Neural Computability and Limits. Theoretical characterizations of neural network computational power have explored universality and limits Siegelmann (1996); Graves (2017); Hahn (2020). Turing completeness has been shown for RNNs Siegelmann (1996) and for attention-based architectures under certain conditions Pérez et al. (2021), though these results require infinite precision, recurrence, or stack memory. Our work instead focuses on bounded computation, demonstrating that finite-depth feedforward networks exactly simulate bounded-memory probabilistic machines.
Latent Space Compression and Geometry. Multiple works have connected neural embeddings with symbolic or geometric structure Hupkes et al. (2020), yet these are largely observational. Our results formally prove that PFA states can be embedded into continuous vector spaces with provable separation and logarithmic compression—extending NFA results from Dhayalkar (2025d) to the probabilistic domain. In particular, we show that probabilistic belief vectors, computed via matrix iteration, can replace explicit path enumeration while maintaining semantic equivalence-offering a differentiable relaxation of PFA inference.
Symbolic Learnability in Neural Models. The supervised learning of regular languages using neural networks has been explored in prior work Butoi et al. (2025), typically treating networks as black-box function approximators. These approaches focus on empirical accuracy while offering limited insight into the underlying symbolic structure or convergence behavior. In contrast, our framework enables direct learning of probabilistic finite automata through symbolic feedforward architectures, where each layer explicitly encodes a stochastic state update. We show that when trained with labeled data using gradient descent, our symbolic PFA simulator converges to an exact reproduction of the underlying probabilistic machine. As established in Proposition 5.1, the trained network not only achieves perfect classification accuracy, but also preserves semantic fidelity to the PFA’s marginal transition structure. This demonstrates that neural networks can serve not merely as approximators, but as exact and interpretable simulators of probabilistic state machines—offering a principled alternative to purely empirical sequence learners.
Summary. In contrast to prior heuristic, empirical, or probing-based efforts, our work presents a complete theory for probabilistic state machine simulation in neural networks—covering belief updates, $\varepsilon$-closures, subset construction, path tracing, and learnability. Our results bridge automata theory and differentiable programming, providing a robust foundation for future work on symbolic reasoning, neural formal languages, and the limits of bounded computation in feedforward systems.
@@ 3 Preliminaries and Formal Definitions
We now formally define the mathematical concepts underlying our framework for simulating probabilistic finite automata (PFAs) using feedforward neural networks. These definitions extend the symbolic constructions from nondeterministic automata defined in  Dhayalkar (2025d) to the probabilistic setting.
@@
NFAs and DFAs are special cases of PFAs where transition matrices are binary-valued ($\{0,1\}$), and the state vector is always a one-hot indicator, as stated in  Dhayalkar (2025d). The PFA formulation generalizes this by allowing mixtures over next states and enabling real-valued, probabilistic computation.
@@ 4 Theoretical Framework
We now present the main theoretical framework for simulating probabilistic finite automata (PFAs) using feedforward neural networks. In contrast to the binary and thresholded updates used for simulating NFAs, PFAs propagate soft state vectors using row-stochastic matrix multiplications. This section establishes the core mathematical foundation for symbolic, differentiable simulation of PFAs.
@@
Proof. The complete proof is provided in Appendix A.1.
@@
Proof. The complete proof is provided in Appendix A.2.
@@
Relation to Prior Work: The classical subset construction for NFAs uses binary vectors to track reachable states under symbol Giles et al. (1992a); Weiss et al. (2018); Dhayalkar (2025d). Here, we replace Boolean logic with stochastic linear operators, resulting in an exact probabilistic analog. This connects symbolic automata theory with probabilistic state-space models and forward inference in stochastic processes, while remaining fully differentiable and compositional.
@@
$\bm{p}^{(k+1)}=\sigma\!\left(\bm{p}^{(k)}T^{\varepsilon}\right),\quad\bm{p}^{(0)}=\bm{p}_{t},$
@@
Proof. The complete proof is provided in Appendix A.3.
@@
Unlike binary $\varepsilon$-closures that track only state reachability, this formulation captures the probability of reaching each state through any number of $\varepsilon$-transitions. The clipping function $\sigma$ ensures convergence by capping activation probabilities at $1$, preventing runaway amplification. This enables neural PFA simulations to perform $\varepsilon$-closure via fully differentiable, parallel matrix-vector iterations, offering a principled and convergent method for modeling probabilistic nondeterminism within a linear algebraic framework.
@@
Let $\mathcal{A}=(Q,\Sigma,\delta,\pi_{0},F)$ be a probabilistic finite automaton (PFA) with $|Q|=n$ states, and let each $T^{x}\in\mathcal{S}_{n}$ denote the stochastic transition matrix for symbol $x\in\Sigma$. Then there exists a feedforward network $f_{\theta}$ that exactly simulates the acceptance probability of $\mathcal{A}$ on any input string $x=x_{1}x_{2}\cdots x_{L}\in\Sigma^{*}$. That is,
$f_{\theta}(x)=\langle\pi_{0}\,T^{x_{1}}T^{x_{2}}\cdots T^{x_{L}},\mathbf{1}_{F}\rangle\;=\;\Pr_{\mathcal{A}}(x).$
In particular, for any $x=x_{1}\cdots x_{L}$, the computation $s_{L}=\pi_{0}T^{x_{1}}\cdots T^{x_{L}}$ induces a feedforward computation graph of depth $L$ (one layer per symbol), with symbol-conditioned transition operators $\{T^{x}\}_{x\in\Sigma}$ shared across layers.
Proof. The complete proof is provided in Appendix A.4.
Explanation: This proposition formalizes the core claim of this paper: feedforward neural networks with shared symbolic parameters can exactly simulate the behavior of probabilistic finite automata over arbitrary-length input sequences. Each layer corresponds to a single input symbol, and matrix multiplication with a stochastic transition matrix propagates the probabilistic state forward.
The network simulates probabilistic automata using soft transitions, maintaining and updating distributions over states rather than binary activations. This aligns with the semantics of forward inference in stochastic processes, allowing the network to compute the likelihood of reaching each final state. The result establishes that any PFA can be exactly represented by a symbolic feedforward network—without recurrence, memory, or sampling—where each layer corresponds to one automaton step. This yields an interpretable, parallelizable architecture that also supports automatic differentiation and gradient-based training of symbolic machines.
Relation to Prior Work: Prior work has used recurrent architectures to simulate automata, often requiring sampling, memory, or training with surrogate losses Weiss et al. (2018); Graves (2017); Hupkes et al. (2020). Our construction shows that even probabilistic automata can be represented in a purely feedforward form. This generalizes recent work on NFA simulation via binary vectors Dhayalkar (2025d) to probabilistic computation, enabling a broader class of symbolic machines to be embedded into deep learning models.

### v2

@@ Abstract
We present a formal and constructive theory showing that probabilistic finite automata can be exactly simulated using symbolic feedforward neural networks. Our architecture represents state distributions as vectors and transitions as stochastic matrices, enabling probabilistic state propagation via matrix-vector products. This yields a parallel, interpretable, and differentiable simulation of probabilistic finite automata dynamics using soft updates without recurrence. We formally characterize probabilistic subset construction, finite $\varepsilon$-closure, and exact simulation via layered symbolic computation, and prove equivalence between probabilistic finite automata and specific classes of neural networks. We further show that these symbolic simulators are not only expressive but learnable: trained with standard gradient descent-based optimization on labeled sequence data, they recover the exact behavior of ground-truth probabilistic finite automata. This learnability, formalized in Proposition 5.1, is the crux of this work. Our results unify probabilistic automata theory with classical neural architectures under a rigorous algebraic framework, bridging the gap between symbolic computation and deep learning.
footnotetext: This paper has been submitted to IEEE Access and is currently under review.
@@ 1 Introduction
This work closes that gap. We show that feedforward networks—without recurrence, memory, or sampling—can simulate PFAs exactly. Each layer in our architecture implements a stochastic state update using matrix-vector multiplication with symbolic transition matrices, followed by a softmax or sigmoid operation for normalization and classification. Our framework extends prior work on DFA Dhayalkar (2025a) and NFA Dhayalkar (2025b) simulation by unifying deterministic, nondeterministic, and probabilistic transitions under a single symbolic neural model.
Formally, we define the simulation of PFAs by feedforward networks as the exact computation of the acceptance probability for any string via symbolic matrix composition. Proposition 4.7 and Theorem 4.10 prove a bidirectional equivalence: (1) every PFA can be simulated by a depth-unrolled feedforward network with shared stochastic parameters, and (2) every such symbolic network corresponds to a valid PFA. This allows us to replace automaton-based reasoning with static, interpretable network computation. Moreover, we define new symbolic mechanisms for probabilistic $\varepsilon$-closure (Lemma 4.6), probabilistic subset construction (Theorem 4.5), and belief state propagation (Proposition 4.1), all implemented as differentiable neural layers.
@@
  • We present the first constructive theory for simulating probabilistic finite automata (PFAs) with $\varepsilon$-transitions using symbolic feedforward networks.
  • We introduce exact formulations of probabilistic subset construction and finite $\varepsilon$-closure via matrix-vector dynamics, enabling differentiable automaton emulation.
@@
This work advances the field of neural-symbolic computation by providing a unified, interpretable, and learnable framework for simulating probabilistic automata in neural networks. It lays the groundwork for deeper integration of formal language theory and deep learning, with applications in structured reasoning, interpretable sequence modeling, and symbolic control.
@@ 2 Related Work
The study of neural networks as computational models has a long and rich history, spanning universal approximation theory Cybenko (1989); Hornik et al. (1989), neural-symbolic integration Besold et al. (2017); d’Avila Garcez et al. (2019), automata learning Weiss et al. (2018); Hupkes et al. (2020); Michalenko (2019), and formal language recognition Hopcroft et al. (2006); Hays (2003). This paper builds on this foundational work by developing a precise theoretical and empirical framework that characterizes how feedforward networks simulate probabilistic finite automata (PFAs), extending prior results from the deterministic setting Dhayalkar (2025a) and non-deterministic setting Dhayalkar (2025b).
2.1 Neural-Symbolic Computation
@@
Early research in neural-symbolic computation sought to integrate symbolic reasoning with neural networks via hybrid architectures or recurrent logic units d’Avila Garcez et al. (2019); Besold et al. (2017). These models demonstrated that certain forms of propositional logic could be encoded and manipulated within connectionist systems, yet often lacked rigorous formalism and focused primarily on logic inference rather than state-machine emulation. More recent efforts have introduced differentiable models of symbolic processes Graves (2017), neural theorem proving Rocktäschel and Riedel (2017), and neural architecture search over structured domains Łukasz Kaiser et al. (2017). However, these methods have been largely heuristic or data-driven, often requiring supervision or pre-specification of symbolic structure.
2.2 Neural Networks as Automata
@@
Multiple works have explored whether and how neural networks implicitly implement automata when trained on regular or formal languages Weiss et al. (2018); Hupkes et al. (2020); Rabinovich et al. (2017). These studies generally relied on probing internal representations of recurrent networks (RNNs, LSTMs, Transformers) to detect automaton-like behaviors such as state memorization or finite counting. Yet, such work remained empirical in nature, offering limited constructive insight into network design or simulation guarantees. Our work, building upon Dhayalkar (2025b), replaces this probing-based approach with a complete, formal construction that shows how feedforward networks can simulate finite automata exactly—transition by transition—both in deterministic and probabilistic settings. We go beyond approximate modeling to provide exact emulation and generalization guarantees.
2.3 Formal Automata Theory
@@
The classical theory of automata Hopcroft et al. (2006); Hays (2003) distinguishes deterministic, nondeterministic, and probabilistic automata (DFAs, NFAs, PFAs), with growing expressive power but also increasing simulation complexity. In the deterministic case, transitions are deterministic functions, while in PFAs, transitions are stochastic matrices operating on probability distributions over states. While PFAs have been studied extensively in formal language theory Rabin (1963); Paz (1971), their neural simulation has received little attention. Most existing neural-symbolic models either avoid stochastic transitions or simulate them via recurrent dynamics, often without exactness guarantees. We fill this gap by providing the first constructive simulation of PFAs using feedforward neural networks, showing that symbolic probabilistic state updates, $\varepsilon$-closures, and subset constructions can all be implemented using static, differentiable architectures.
2.4 Probabilistic and Nondeterministic Extensions
@@
Recent advances in learning finite automata machines Siegelmann (1996); Giles et al. (1992b); Hong et al. (2022); Omlin and Giles (1996); Marzen and Crutchfield (2022); Wei et al. (2022) explore how RNNs approximate NFAs and pushdown automata, yet lack constructive bounds or clarity about how transitions or path probabilities are represented. Prior work by Dhayalkar (2025a) introduced a constructive theory for simulating deterministic finite automata (DFAs) using depth-unrolled feedforward neural networks with ReLU or threshold activations. That work demonstrated exact transition emulation, formally proved state transition separability and embedding capacity, and connected network representations to the classical Myhill-Nerode theorem. Subsequently, this framework was extended to the nondeterministic case Dhayalkar (2025b), showing that feedforward architectures could simulate nondeterministic finite automata (NFAs) without requiring explicit path enumeration. The NFA variant introduced symbolic binary state vectors and unrolled subset constructions to represent all reachable states at each timestep. Importantly, it preserved full theoretical interpretability while maintaining exact correspondence with NFA behavior, including support for $\varepsilon$-transitions.
@@
2.5 Neural Computability and Limits
Theoretical characterizations of neural network computational power have explored universality and limits Siegelmann (1996); Graves (2017); Hahn (2020). Turing completeness has been shown for RNNs Siegelmann (1996) and for attention-based architectures under certain conditions Pérez et al. (2021), though these results require infinite precision, recurrence, or stack memory. Our work instead focuses on bounded computation, demonstrating that finite-depth feedforward networks exactly simulate bounded-memory probabilistic machines.
2.6 Latent Space Compression and Geometry
Multiple works have connected neural embeddings with symbolic or geometric structure Hupkes et al. (2020), yet these are largely observational. Our results formally prove that PFA states can be embedded into continuous vector spaces with provable separation and logarithmic compression—extending NFA results from Dhayalkar (2025b) to the probabilistic domain. In particular, we show that probabilistic belief vectors, computed via matrix iteration, can replace explicit path enumeration while maintaining semantic equivalence-offering a differentiable relaxation of PFA inference.
2.7 Symbolic Learnability in Neural Models
The supervised learning of regular languages using neural networks has been explored in prior work Butoi et al. (2025), typically treating networks as black-box function approximators. These approaches focus on empirical accuracy while offering limited insight into the underlying symbolic structure or convergence behavior. In contrast, our framework enables direct learning of probabilistic finite automata through symbolic feedforward architectures, where each layer explicitly encodes a stochastic state update. We show that when trained with labeled data using gradient descent, our symbolic PFA simulator converges to an exact reproduction of the underlying probabilistic machine. As established in Proposition 5.1, the trained network not only achieves perfect classification accuracy, but also preserves semantic fidelity to the PFA’s marginal transition structure. This demonstrates that neural networks can serve not merely as approximators, but as exact and interpretable simulators of probabilistic state machines—offering a principled alternative to purely empirical sequence learners.
2.8 Summary
In contrast to prior heuristic, empirical, or probing-based efforts, our work presents a complete theory for probabilistic state machine simulation in neural networks—covering belief updates, $\varepsilon$-closures, subset construction, path tracing, and learnability. Our results bridge automata theory and differentiable programming, providing a robust foundation for future work on symbolic reasoning, neural formal languages, and the limits of bounded computation in feedforward systems.
@@ 3 Preliminaries and Formal Definitions
We now formally define the mathematical concepts underlying our framework for simulating probabilistic finite automata (PFAs) using feedforward neural networks. These definitions extend the symbolic constructions from nondeterministic automata defined in Dhayalkar (2025b) to the probabilistic setting.
@@
Definition 3.4 ($\varepsilon$-Closure).
Given a state $q\in Q$, the $\varepsilon$-closure of $q$, denoted $\mathrm{cl}_{\varepsilon}(q)$, is the set of states reachable from $q$ via a sequence of zero or more $\varepsilon$-transitions. For a set of states $S\subseteq Q$, we define: $\mathrm{cl}_{\varepsilon}(S)=\bigcup_{q\in S}\mathrm{cl}_{\varepsilon}(q).$
@@
NFAs and DFAs are special cases of PFAs where transition matrices are binary-valued ($\{0,1\}$), and the state vector is always a one-hot indicator, as stated in Dhayalkar (2025a; b). The PFA formulation generalizes this by allowing mixtures over next states and enabling real-valued, probabilistic computation.
@@ 4 Theoretical Framework
We now present the main theoretical framework for simulating probabilistic finite automata (PFAs) using feedforward neural networks. In contrast to the binary and thresholded updates used for simulating NFAs Dhayalkar (2025b), PFAs propagate soft state vectors using row-stochastic matrix multiplications. This section establishes the core mathematical foundation for symbolic, differentiable simulation of PFAs.
@@
Proof Sketch. The proof sketch is provided in Appendix A.1.
@@
Proof Sketch. The proof sketch is provided in Appendix A.2.
@@
Relation to Prior Work: The classical subset construction for NFAs uses binary vectors to track reachable states under symbol Giles et al. (1992a); Weiss et al. (2018); Dhayalkar (2025b). Here, we replace Boolean logic with stochastic linear operators, resulting in an exact probabilistic analog. This connects symbolic automata theory with probabilistic state-space models and forward inference in stochastic processes, while remaining fully differentiable and compositional.
@@
$\bm{p}^{(k+1)}=\bm{p}^{(k)}T^{\varepsilon},\quad\bm{p}^{(0)}=\bm{p}_{t},$
@@
Proof Sketch. The proof sketch is provided in Appendix A.3.
@@
Unlike binary $\varepsilon$-closures that track only state reachability, this formulation captures the probability of reaching each state through any number of $\varepsilon$-transitions. Because $T^{\varepsilon}$ is row-substochastic, each update preserves normalization and ensures boundedness, while monotonicity guarantees convergence. This enables neural PFA simulations to perform $\varepsilon$-closure via fully differentiable, parallel matrix-vector iterations, offering a principled and convergent method for modeling probabilistic nondeterminism within a linear algebraic framework.
@@
Let $\mathcal{A}=(Q,\Sigma\cup\{\varepsilon\},\delta,\pi_{0},F)$ be a probabilistic finite automaton (PFA) with $|Q|=n$ states, where each $T^{x}\in\mathcal{S}_{n}$ denotes the stochastic transition matrix for symbol $x\in\Sigma$, and $T^{\varepsilon}\in\mathcal{S}_{n}$ denotes the stochastic $\varepsilon$-transition matrix. Then there exists a feedforward network $f_{\theta}$ that exactly simulates the acceptance probability of $\mathcal{A}$ on any input string $x=x_{1}x_{2}\cdots x_{L}\in\Sigma^{*}$. That is,
$f_{\theta}(x)=\Big\langle\pi_{0}\Big(\sum_{m=0}^{n-1}(T^{\varepsilon})^{m}\Big)T^{x_{1}}\Big(\sum_{m=0}^{n-1}(T^{\varepsilon})^{m}\Big)\cdots\\ T^{x_{L}}\Big(\sum_{m=0}^{n-1}(T^{\varepsilon})^{m}\Big),\mathbf{1}_{F}\Big\rangle\;=\;\Pr_{\mathcal{A}}(x).$ (1)
In particular, for any $x=x_{1}\cdots x_{L}$, the computation
$s_{L}=\pi_{0}\Big(\sum_{m=0}^{n-1}(T^{\varepsilon})^{m}\Big)T^{x_{1}}\Big(\sum_{m=0}^{n-1}(T^{\varepsilon})^{m}\Big)\cdots T^{x_{L}}\Big(\sum_{m=0}^{n-1}(T^{\varepsilon})^{m}\Big)$
induces a feedforward computation graph of depth $L$ (one layer per symbol), with symbol-conditioned transition operators $\{T^{x}\}_{x\in\Sigma}$ shared across layers.
Proof Sketch. The proof sketch is provided in Appendix A.4.
Explanation: This proposition formalizes the core claim of this paper: feedforward neural networks with shared symbolic parameters can exactly simulate the behavior of probabilistic finite automata with $\varepsilon$-transitions. Each layer corresponds to a single input symbol, and matrix multiplication with the corresponding transition matrix propagates the probabilistic state forward, while the $\varepsilon$-closure is computed as mentioned in Lemma 4.6. The network thus simulates probabilistic automata using soft transitions, maintaining distributions over states rather than binary activations, while exactly incorporating $\varepsilon$-steps. This yields an interpretable architecture that also supports automatic differentiation and gradient-based training of symbolic machines.
Relation to Prior Work: Prior work has used recurrent architectures to simulate automata, often requiring sampling, memory, or training with surrogate losses Weiss et al. (2018); Graves (2017); Hupkes et al. (2020). Our construction shows that even probabilistic automata can be represented in a purely feedforward form. This generalizes recent work on NFA simulation via binary vectors Dhayalkar (2025b) to probabilistic computation, enabling a broader class of symbolic machines to be embedded into deep learning models.